OpenStudy (anonymous):

Find (A ◦ r)(x). A = π r^2 r =x/2

OpenStudy (anonymous):

where you see an \(r\) replace it by \(\frac{x}{2}\)

OpenStudy (anonymous):

a(r(x)) so π(x/2)^2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1. A = π x 2. None of these 3. A = 2π x 4. A = 1 4 π x2 5. A = 4π x2 6. A = 1 2 π x 7. A = π x2

OpenStudy (anonymous):

which one then

OpenStudy (anonymous):

1. A = π x 2. None of these 3. A = 2π x 4. A =1/4π x^2 5. A = 4π x^2 6. A =1/2π x 7. A = π x2 Sorry about that

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

\[\frac{\pi x^2}{4}\] if that is a choice

OpenStudy (anonymous):

divided by 2